OpenStudy (anonymous):

Factor completely 4x2 + 25x + 6. (4x + 1)(x + 6) (4x + 6)(x + 1) (2x + 3)(2x + 2) (2x + 6)(2x + 1)

OpenStudy (anonymous):

sorry it is 4x^2

OpenStudy (nurali):

4x^2 + 25x + 6 = 4x^2 + 24x + x + 6 = 4x(x + 6) + 1(x + 6) = (x + 6)(4x + 1)

OpenStudy (anonymous):

Thank you so much!! That makes a lot of sense!

OpenStudy (nurali):

Anytime.